OpenStudy (anonymous):

How do you solve this problem 6+2i / 5i ?

OpenStudy (amistre64):

solve means different things in different context; what is it you are spose to do for a solution?

OpenStudy (amistre64):

we can just plain divide, but thats not really a 'solve' thing

OpenStudy (amistre64):

we can multiply top and bottom by 5i; but thats more simplifiaction than a solve

OpenStudy (amistre64):

\[\frac{6+2i}{5i}*\frac{5i}{5i}=\frac{30i -10}{-25}\] \[-{6i -5\over 5}\]